OpenStudy (anonymous):

Find the angles O between -180 and 180 for which cos ) = 0.940.

OpenStudy (maheshmeghwal9):

\[\LARGE{\color{red}{0.940\approx 1=\cos 0^o=\cos 0*\pi.}}\]